# ISDrama: Immersive Spatial Drama Generation through Multimodal Prompting

### v1

@@ 4.3 Immersive Drama Transfomer
Drama-MOE. To enhance prosodic expressiveness and pose control, we propose Drama-MOE (Mixture of Experts), selecting suitable experts based on various input conditions. As shown in Figure 4 (d), Drama-MOE consists of two expert groups, each focusing on dramatic prosody and spatial information. Prosodic MOE leverages prosody in aligned prompt audio embedding $z_{a}$ and semantics in content embedding $z_{c}$ to select suitable experts for fine-grained dramatic prosody modeling, such as an expert specialized in a happy, fast, high-pitched tone. $z_{a}$ is aligned with the inputs by a cross-attention model. The Spatial MOE conditions on $z_{a}$, adjusting inputs to match the corresponding spatial information, like sound source pose-induced changes and binaural differences in phase and loudness. It selects experts based on inputs, like one specialized in a sound source slowly approaching the listener from the far left direction.
@@
$\phi(x)\triangleq[\cos(W_{p}x)||\sin(W_{p}x)||\sigma(B_{\bar{p}}+W_{\bar{p}}x)],$ (3)
where $W_{p},W_{\bar{p}}$, and $B_{\bar{p}}$ are learnable parameters. The output of FAN combines sinusoidal functions and a nonlinear activation, capturing periodic and non-periodic variations. This enhances the expressive power for spatial-temporal and time-frequency auditory modeling, aiding in the synchronized modeling of speech prosody and spatial information.
@@
$\mathcal{L}_{balance}=\alpha N\sum_{i=1}^{N}\left(\frac{1}{B}\sum_{h\in B}g(h)_{i}\right),$ (4)
where $B$ is the batch size, $N$ is the number of experts, and $\alpha$ a hyperparameter controlling regularization strength. For more details, please refer to Appendix C.7.
4.4 Complete Drama Inference Procedure
@@
$v_{cfg}(x,t|z_{p};\theta)=\gamma\alpha v_{t}(x,t|a,C;\theta)+$ (5)
@@
where $\gamma$ is the CFG scale that balances creativity and controllability, while $\alpha$ balances context consistency and input controllability. Setting $\gamma=3$ and $\alpha=0.4$, we improve generation quality and incorporate previously generated audio to enhance the prosody consistency of the same speaker within a drama act. This ensures coherence while preserving the timbre and accent of the original prompt audio.
Moreover, as prosody can be learned from previous prompt audio in the same context, this approach also improves the expressiveness of semantically aligned prosody. Additionally, by leveraging the accelerated inference capabilities of flow matching and Mamba blocks, our model can efficiently generate immersive spatial drama. For more details, please refer to Appendix C.2 and C.3.
5 Experiments
5.1 Experimental Setup
Implementation Details. Mel-spectrograms are derived from raw binaural waveforms with a 48 kHz sample rate, 1024 window size, 256 hop size, and 80 mel bins. We use four Mamba-Transformer blocks. The flow-matching timestep is 1000 for training and 25 for inference with the Euler ODE solver. For the training procedure of Immersive Drama Transformer, we use 8 NVIDIA RTX-4090 GPUs with a batch size of 12K frames per GPU for 100K steps. The Adam optimizer is applied with a learning rate of $5\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and 10K warm-up steps. Please refer to Appendix C.1 for more details.
Evaluation Metrics. We perform both subjective and objective evaluations on the generated samples. To ensure a fair comparison with existing monaural speech synthesis models, we adopt monaural evaluation metrics and then evaluate binaural metrics for cascade-generated binaural speech with BinauralGrad (Leng et al., 2022). 1) For objective evaluation, we use Character Error Rate (CER) and Cosine Similarity (SIM) to assess the content accuracy and speaker similarity with prompt audio. F0 Frame Error (FFE) is used to evaluate the quality of prosody modeling. Since existing binaural metrics are scarce and not suitable for our one-stage binaural speech generation task, we have designed several new metrics. We extract Interaural Phase Difference (IPD) and Interaural Level Difference (ILD) from the binaural mel-spectrograms and compute MAE with GT. We also compute cosine similarity of angle and distance embedding extracted from SPATIAL-AST (Zheng et al., 2024) with GT for spatial evaluation. 2) For subjective evaluation, we conduct Mean Opinion Score (MOS), which is rated from 1 to 5 and reported with 95% confidence intervals. MOS-Q evaluates the synthesized quality (like naturalness, spatial perception, and coherence), MOS-S assesses speaker similarity in timbre and accent, and MOS-E measures the expressiveness of semantically aligned prosody. For binaural metrics, we employ MOS-P to evaluate the pose consistency between the multimodal prompt and the generated audio. In the ablation study, we conduct Comparative Mean Opinion Score (CMOS). Please refer to Appendix D for more details,
Table 2: Monaural speech quality comparison. For testing quality and speaker similarity, we use single-sentence speech for evaluation.
@@ 5.1 Experimental Setup
Table 3: Binaural speech quality comparison. We use complete drama for spatial evaluation. ANG and Dis denote angle and distance. Spatialization refers to the generation of binaural audio directly from the GT monaural audio based on the geometric pose.
@@
5.2 Results
Comparison of baseline models for monaural speech. To ensure a fair evaluation of the synthesized quality, speaker similarity to the prompt audio, and prosodic expressiveness, we compute monaural speech metrics after averaging the binaural speech generated by ISDrama across channels. For the baseline models, we employ several strong popular open-source speech synthesis models, including: 1) Uniaudio (Yang et al., 2023), 2) StyleTTS 2 (Li et al., 2024c), 3) CosyVoice (Du et al., 2024), 4) FireRedTTS (Guo et al., 2024), and 5) F5-TTS (Chen et al., 2024). We use their open-source codes on GitHub and train them using our MRSDrama dataset in monaural speech format. We also provide a detailed introduction for each baseline model in Appendix E.1. To achieve a more precise evaluation of quality and speaker similarity, we conduct the assessment using only single sentences for all models. Table 2 shows that ISDrama performs well across all monaural metrics. In synthesized quality (CER, MOS-Q) and speaker similarity (SIM, MOS-S), it is comparable to the performance of the best speech synthesis baseline model, while it surpasses all baseline models in prosodic expressiveness (FFE, MOS-E). These well-performing results can be attributed to the integration of semantics to model dramatic prosody with semantically aligned emotion and rhythm in Drama-MOE, as well as F0 supervision during training.
Comparison of Baseline Models for Binaural Speech. Based on the subjective and objective spatial metrics we designed, we train a commonly used monaural-to-binaural spatialization model, BinauralGrad (Leng et al., 2022), on our MRSDrama dataset to convert the generated monaural speech from baseline models, conditioned on geometric poses, into binaural speech for evaluating spatial performance. To better test the coherence of spatial information and the consistency of prosody, we concatenate sentences from the same generated drama for evaluation. As shown in Table 3, with the same geometric pose input, the binaural speech generated by ISDrama outperforms all two-stage models across all spatial metrics, both subjective and objective. The high MOS-Q shows coherence in continuous multi-speaker speech generation, which can be attributed to the effectiveness of our context-consistent CFG in generating complete drama. In more flexible application scenarios, including video and textual prompt inputs, ISDrama continues to deliver excellent performance, thanks to the unified pose representation extracted by the Multimodal Pose Encoder. Although the textual prompt leads to lower objective metrics, it achieves a strong MOS-P score, which can be attributed to the general nature of the text descriptions. Specifically, the same textual description can generate multiple different pose sequences, resulting in spatial information not fully aligned with the ground truth.
Table 4: Ablation studies on Immersive Drama Transformer. MPE denotes Multimodal Pose Encoder.
@@ 5.2 Results
When listening to the demos and considering the pose cues provided by multimodal inputs, it is evident that ISDrama can maintain consistency with the poses of the sound source and demonstrate coherent and expressive prosody aligned with semantics. Overall, ISDrama effectively models spatial information and prosodic expressiveness, achieving a well-balanced outcome. These results can be attributed to the exceptional pose modeling capabilities of our Multimodal Pose Encoder, as well as the enhanced dramatic prosody and pose control facilitated by Immersive Drama Transformer.
5.3 Ablation Study
Multimodal Pose Encoder. Tables 3 and 4 present the results of feeding pose embeddings of different prompts into the Immersive Drama Transformer, with or without contrastive learning. We observe that the absence of contrastive learning has minimal impact on geometric poses. This is because the geometric pose embeddings include 3D orientation, quaternion orientation, and radial velocity, which provide sufficient information for accurate pose modeling. However, for silent video and textual prompts, omitting contrastive learning significantly reduces spatial performance. This highlights the importance of contrastive learning, as it enhances the model’s ability to generate unified embeddings from multimodal prompts, which is crucial for supporting diverse and flexible application scenarios.
@@ 5.3 Ablation Study
Immersive Drama Transformer. We evaluate the effects of removing the supervision of F0, replacing Mamba with a memory-equivalent Transformer, and eliminating the context-consistent CFG. The results are shown in Table 4. It can be observed that F0 prediction is crucial for accurate prosody modeling, as its absence results in a significant loss of prosodic quality. The Mamba block, being lightweight, allows stacking more layers than a traditional Transformer with the same memory budget, which leads to further improvements in quality. Additionally, removing the context-consistent CFG reduces the consistency and prosodic expressiveness of the generated outputs. These findings underscore the effectiveness of these design choices in enhancing quality and prosody modeling capabilities of the Immersive Drama Transformer. For more extended results, please refer to Appendix E.
6 Conclusion
In this paper, we introduce a novel task Multimodal Immersive Spatial Drama Generation, focusing on creating continuous multi-speaker binaural speech with dramatic prosody based on multimodal prompts. To support this task, we present MRSDrama, the first multimodal recorded spatial drama dataset, comprising binaural drama audios, scripts, videos, geometric poses, and textual prompts. Then, we propose ISDrama, the first immersive spatial drama generation model based on multimodal prompting. To extract a unified pose representation from multimodal prompts, we design the Multimodal Pose Encoder, a contrastive learning-based framework, accounting for the Doppler effect caused by moving speakers. To generate immersive spatial drama effectively and stably, we develop the Immersive Drama Transformer, a flow-based Mamba-Transformer model, incorporating Drama-MOE, which selects proper experts to enhance prosodic expressiveness and pose control. Then, we adopt a context-consistent CFG strategy to coherently generate high-quality complete drama. Experimental results show that ISDrama achieves better performance than baseline models across both objective and subjective metrics.
Impact Statement
Large-scale multimodal datasets and generative models often pose ethical challenges. Since MRSDrama includes human voice and video data, we have implemented data desensitization measures. Specifically, we applied masks to the faces in videos to prevent the leakage of biometric information. Additionally, ISDrama, with its immersive spatial drama generation capabilities, could be misused for applications like dubbing in entertainment short videos, potentially infringing on the copyrights of well-known actors. Furthermore, its ability to transfer and control prosody and spatial poses lowers the technical barriers to generating high-quality drama, which could lead to risks such as unfair competition and potential unemployment for professionals in related theatrical occupations. To mitigate these risks, we plan to explore methods such as audio watermarking to safeguard individual privacy and address misuse concerns.
@@ A.1 Doppler Effect
$f_{o}=\left(\frac{c}{c\pm v_{rad}}\right)f_{s},$ (6)
@@
A.2 Rectified flow-matching
In generative modeling, the true data distribution is denoted as $q(x_{1})$, which can be sampled but lacks an accessible density function. Consider a probability path $p_{t}(x_{t})$, where $x_{0}\sim p_{0}(x)$ represents a standard Gaussian, and $x_{1}\sim p_{1}(x)$ approximates the real data distribution. The core of the flow-matching approach (Liu et al., 2022b) is to model this path directly, which is governed by the following ordinary differential equation (ODE):
$\mathrm{d}x=u(x,t)\mathrm{d}t,\quad t\in[0,1],$ (7)
@@ A.2 Rectified flow-matching
$\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t,p_{t}(x)}\left\|v(x,t;\theta)-u(x,t)\right\|^{2},$ (8)
@@
$\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,p_{1}(x_{1}),p_{t}(x|x_{1})}\left\|v(x,t|C;\theta)-u(x,t|x_{1},C)\right\|^{2}.$ (9)
@@
$\left\|v(x,t|C;\theta)-(x_{1}-x_{0})\right\|^{2}.$ (11)
@@
$x_{t+\epsilon}=x+\epsilon v(x,t|C;\theta),$ (12)
@@
A.3 Mamba
Models based on Structured State Space (SSM), such as S4 and Mamba (Gu & Dao, 2023), draw inspiration from continuous systems that map a 1D sequence $x(t)\in\mathbb{R}$ to another sequence $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. In this system, the matrix $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ serves as the state evolution parameter, while $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times\mathtt{N}}$ are used for projection. The continuous system is:
$h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$ (13)

### v6

@@ 4.3. Immersive Drama Transfomer
Drama-MOE. To enhance prosodic expressiveness and pose control, we propose Drama-MOE (Mixture of Experts), which selects suitable experts based on various input conditions. As shown in Figure 4 (d), our Drama-MOE consists of two expert groups, each focusing on dramatic prosody and spatial information. Prosodic MOE leverages prosody in aligned prompt audio embedding $z_{a}$ and semantics in content embedding $z_{c}$ to select suitable experts for fine-grained dramatic prosody modeling, such as an expert specialized in a happy, fast, high-pitched tone. $z_{a}$ is aligned with the inputs by a cross-attention model. The Spatial MOE conditions on $z_{a}$, adjusting inputs to match the corresponding spatial information, such as sound source pose-induced changes and binaural differences in phase and loudness. It selects proper experts for different inputs, like one expert specialized in a sound source slowly approaching the listener from the far left-front direction.
@@
(3) $\phi(x)\triangleq[\cos(W_{p}x)||\sin(W_{p}x)||\sigma(B_{\bar{p}}+W_{\bar{p}}x)],$
where $W_{p},W_{\bar{p}}$, and $B_{\bar{p}}$ are learnable parameters. FAN combines sinusoidal functions and a nonlinear activation, capturing periodic and non-periodic variations. This enhances the power for spatial-temporal and time-frequency auditory modeling, aiding in the synchronized modeling of speech prosody and spatial information.
@@
(4) $\mathcal{L}_{balance}=\alpha N\sum_{i=1}^{N}\left(\frac{1}{B}\sum_{h\in B}g(h)_{i}\right),$
where $B$ is the batch size, $N$ is the number of experts, and $\alpha$ a hyperparameter controlling regularization strength. For more details and the algorithm, please refer to Appendix C.7.
4.4. Complete Drama Inference Procedure
@@
(5) $v_{cfg}(x,t|z_{p};\theta)=\gamma\alpha v_{t}(x,t|a,C;\theta)+$
@@
where $\gamma$ is the CFG scale that balances creativity and controllability, while $\alpha$ balances context consistency and controllability. Setting $\gamma=3$ and $\alpha=0.4$, we improve generation quality and add previously generated audio for the prosody consistency of the same speaker within a single drama act. This ensures coherence while preserving the timbre and accent of the original prompt audio. Moreover, as prosody can be learned from previously generated audio in the same context, this approach also helps the modeling of semantically aligned prosody. For more details, please refer to Appendix C.3.
5. Experiments
Table 2. Monaural speech quality comparison. For testing quality and speaker similarity, we evaluate single-sentence speech.
@@ 5. Experiments
Table 3. Binaural speech quality comparison. We evaluate complete drama. ANG and Dis denote angle and distance. Spatialization refers to the generation of binaural audio directly from the GT monaural audio input based on the geometric pose.
@@
5.1. Experimental Setup
Implementation Details. Mel-spectrograms are derived from raw binaural waveforms with a 48 kHz sample rate, 1024 window size, 256 hop size, and 80 mel bins. We use four Mamba-Transformer blocks, each of which includes three Mamba blocks. The flow-matching timestep is 1000 for training and 25 for inference with the Euler ODE solver. For the training procedure of Immersive Drama Transformer, we use 8 NVIDIA RTX-4090 GPUs with a batch size of 12K frames per GPU for 100K steps. The Adam optimizer is applied with a learning rate of $5\times 10^{-5}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$, and 10K warm-up steps. Please refer to Appendix C.1 for more details.
Evaluation Metrics. We perform both subjective and objective evaluations on the generated samples. To ensure a fair comparison with existing monaural speech synthesis models, we adopt monaural evaluation metrics and then evaluate binaural metrics for cascade-generated binaural speech with BinauralGrad (Leng et al., 2022). 1) For objective evaluation, we use Character Error Rate (CER) and Cosine Similarity (SIM) to assess the content accuracy and speaker similarity with prompt audio. F0 Frame Error (FFE) is used to evaluate the quality of prosody modeling. Since existing binaural metrics are scarce and not suitable for our one-stage binaural speech generation task, we have designed several new metrics. We extract Interaural Phase Difference (IPD) and Interaural Level Difference (ILD) from the binaural mel-spectrograms and compute MAE with GT. We also compute cosine similarity of angle and distance embedding extracted from SPATIAL-AST (Zheng et al., 2024) with GT for spatial evaluation. 2) For subjective evaluation, we conduct Mean Opinion Score (MOS), which is rated from 1 to 5 and reported with 95% confidence intervals. MOS-Q evaluates the synthesized quality (like naturalness, spatial perception, and coherence), MOS-S assesses speaker similarity in timbre and accent, and MOS-E measures the expressiveness of semantically aligned prosody. For binaural metrics, we employ MOS-P to evaluate the pose consistency between the multimodal prompt and the generated audio. In the ablation study, we conduct Comparative Mean Opinion Score (CMOS). For more evaluation details, please refer to Appendix D.
Table 4. Ablation studies on Immersive Drama Transformer. MPE denotes Multimodal Pose Encoder.
@@ 5.1. Experimental Setup
5.2. Results
Comparison of baseline models for monaural speech. To ensure a fair evaluation of the synthesized quality, speaker similarity to the prompt audio, and prosodic expressiveness, we compute monaural speech metrics after averaging the binaural speech generated by ISDrama across channels. For the baseline models, we employ several strong popular open-source speech synthesis models, including: 1) Uniaudio (Yang et al., 2023), 2) StyleTTS 2 (Li et al., 2024a), 3) CosyVoice (Du et al., 2024), 4) FireRedTTS (Guo et al., 2024), and 5) F5-TTS (Chen et al., 2024). We use their open-source codes on GitHub and train them using our MRSDrama dataset in monaural speech format. We also provide a detailed introduction for each baseline model in Appendix E.1. To achieve a more precise evaluation of quality and speaker similarity, we conduct the assessment using only single sentences for all models. Table 2 shows that ISDrama performs well across all monaural metrics. In synthesized quality (CER, MOS-Q) and speaker similarity (SIM, MOS-S), it is comparable to the performance of the best speech synthesis baseline model, while it surpasses all baseline models in prosodic expressiveness (FFE, MOS-E). These well-performing results can be attributed to the integration of semantics to model dramatic prosody with semantically aligned emotion and rhythm in Drama-MOE, as well as F0 supervision during training.
Comparison of Baseline Models for Binaural Speech. Based on the subjective and objective spatial metrics we designed, we train a commonly used monaural-to-binaural spatialization model, BinauralGrad (Leng et al., 2022), on our MRSDrama dataset to convert the generated monaural speech from baseline models, conditioned on geometric poses, into binaural speech for evaluating spatial performance. To better test the coherence of spatial information and the consistency of prosody, we concatenate sentences from the same generated drama for evaluation. As shown in Table 3, with the same geometric pose input, the binaural speech generated by ISDrama outperforms all two-stage models across all spatial metrics. The high MOS-Q shows coherence in continuous multi-speaker speech generation, which can be attributed to the effectiveness of our context-consistent CFG in generating complete drama. In more flexible scenarios, including video and textual prompt inputs, ISDrama continues to deliver excellent performance, thanks to the unified pose representation extracted by the Multimodal Pose Encoder. Although the textual prompt leads to lower objective metrics, it achieves a strong MOS-P score, which can be attributed to the general nature of the text descriptions. Specifically, the same textual description can generate multiple pose sequences, resulting in spatial information not fully aligned with the ground truth.
5.3. Ablation Study
Multimodal Pose Encoder. Tables 3 and 4 present the results of feeding pose embeddings of different prompts into the Immersive Drama Transformer, with or without contrastive learning. We observe that the absence of contrastive learning has minimal impact on geometric poses. This is because the geometric pose embeddings include 3D orientation, quaternion orientation, and radial velocity, which provide sufficient information for accurate pose modeling. However, for silent video and textual prompts, omitting contrastive learning significantly reduces spatial performance. This highlights the importance of contrastive learning, as it enhances the model’s ability to generate unified embeddings from multimodal prompts, which is crucial for supporting a wide range of diverse and flexible immersive spatial drama application scenarios.
@@ 5.3. Ablation Study
Immersive Drama Transformer. We evaluate the effects of removing the supervision of F0, replacing Mamba with a memory-equivalent Transformer, and eliminating the context-consistent CFG. The results are shown in Table 4. It can be observed that F0 prediction is crucial for accurate prosody modeling, as its absence results in a significant loss of prosodic quality. The Mamba block, being lightweight, allows stacking more layers than a traditional Transformer with the same memory budget, which leads to further improvements in quality. Additionally, removing the context-consistent CFG reduces the consistency and prosodic expressiveness of the generated outputs. These findings underscore the effectiveness of these design choices in enhancing quality and prosody modeling capabilities of the Immersive Drama Transformer.
6. Conclusion
In this paper, we introduce a novel task, Multimodal Immersive Spatial Drama Generation, focusing on creating continuous multi-speaker binaural speech with dramatic prosody based on multimodal prompts. To support this task, we present MRSDrama, the first multimodal recorded spatial drama dataset, comprising binaural drama audios, scripts, videos, geometric poses, and textual prompts. Then, we propose ISDrama, the first immersive spatial drama generation model based on multimodal prompting. To extract a unified pose representation from multimodal prompts, we design the Multimodal Pose Encoder, a contrastive learning-based framework. To generate immersive spatial drama effectively and stably, we develop the Immersive Drama Transformer, a flow-based Mamba-Transformer model, incorporating Drama-MOE, which selects proper experts to enhance prosodic expressiveness and pose control. Then, we design a context-consistent CFG strategy to coherently generate complete drama. Experimental results show that ISDrama achieves better performance than baseline models.
Acknowledgements
This work was supported by the National Key R&D Program of China (2022ZD0162000) and the National Natural Science Foundation of China under Grant No.U24A20326.
@@ A.1. Doppler Effect
(6) $f_{o}=\left(\frac{c}{c\pm v_{rad}}\right)f_{s},$
@@
A.2. Rectified flow-matching
In generative modeling, the true data distribution is denoted as $q(x_{1})$, which can be sampled but lacks an accessible density function. Consider a probability path $p_{t}(x_{t})$, where $x_{0}\sim p_{0}(x)$ represents a standard Gaussian, and $x_{1}\sim p_{1}(x)$ approximates the real data distribution. The core of the flow-matching approach (Liu et al., 2022a) is to model this path directly, which is governed by the following ordinary differential equation (ODE):
(7) $\mathrm{d}x=u(x,t)\mathrm{d}t,\quad t\in[0,1],$
@@ A.2. Rectified flow-matching
(8) $\mathcal{L}_{\mathrm{FM}}(\theta)=\mathbb{E}_{t,p_{t}(x)}\left\|v(x,t;\theta)-u(x,t)\right\|^{2},$
@@
(9) $\mathcal{L}_{\mathrm{CFM}}(\theta)=\mathbb{E}_{t,p_{1}(x_{1}),p_{t}(x|x_{1})}\left\|v(x,t|C;\theta)-u(x,t|x_{1},C)\right\|^{2}.$
@@
(11) $\left\|v(x,t|C;\theta)-(x_{1}-x_{0})\right\|^{2}.$
@@
(12) $x_{t+\epsilon}=x+\epsilon v(x,t|C;\theta),$
@@
A.3. Mamba
Models based on Structured State Space (SSM), such as S4 and Mamba (Gu and Dao, 2023), draw inspiration from continuous systems that map a 1D sequence $x(t)\in\mathbb{R}$ to another sequence $y(t)\in\mathbb{R}$ through a hidden state $h(t)\in\mathbb{R}^{\mathtt{N}}$. In this system, the matrix $\mathbf{A}\in\mathbb{R}^{\mathtt{N}\times\mathtt{N}}$ serves as the state evolution parameter, while $\mathbf{B}\in\mathbb{R}^{\mathtt{N}\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times\mathtt{N}}$ are used for projection. The continuous system is:
(13) $h^{\prime}(t)$ $=\mathbf{A}h(t)+\mathbf{B}x(t),$